Practice problems 1. Freshman chemistry courses teach the ideal gas law: pV-nRT. The units are: V (m3), n (mol), R (kg.m2/°K-mol-s2) and T (°K), what are the units of the p component?
Answer #1

putting all the units in the ideal gas equation

[p]*m3 = mol* (kg.m2/K.mol.s2) * K

simplifying

[p] = kg/(m.s2)

now the unit of pressure, Pascal = kg/(m.s2)

hence the unit of p component is Pascal